OpenStudy (anonymous):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation?

OpenStudy (anonymous):

which system?

OpenStudy (anonymous):

2x + y = 8 -x - 3y = -12

OpenStudy (anonymous):

-x-3y=-12=>12-3y=x or x=12-3y

OpenStudy (anonymous):

the answers are , 3y + 12 -3y + 12 3y - 12 -3y - 12 3y + 12

OpenStudy (anonymous):

12-3y is the same as -3y+12

OpenStudy (anonymous):

ohh yeahhh truee thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Nick works two jobs to pay for college. He tutors for $15 per hour and also works as a bag boy for $8 per hour. Due to his class and study schedule, Nick is only able to work up to 20 hours per week, but must earn at least $150 per week. If t represents the number of hours Nick tutors and b represents the number of hours he works as a bag boy, which system of inequalities represents this scenario?